# A general approach to modeling environmental mixtures with multivariate outcomes

### v1

@@ Abstract
An important goal of environmental health research is to assess the health risks posed by mixtures of multiple environmental exposures. In these mixtures analyses, flexible models like Bayesian kernel machine regression and multiple index models are appealing because they allow for arbitrary non-linear exposure-outcome relationships. However, this flexibility comes at the cost of low power, particularly when exposures are highly correlated and the health effects are weak, as is typical in environmental health studies. We propose an adaptive index modelling strategy that borrows strength across exposures and outcomes by exploiting similar mixture component weights and exposure-response relationships. In the special case of distributed lag models, in which exposures are measured repeatedly over time, we jointly encourage co-clustering of lag profiles and exposure-response curves to more efficiently identify critical windows of vulnerability and characterize important exposure effects. We then extend the proposed approach to the multivariate index model setting where the true index structure—the number of indices and their composition—is unknown, and introduce variable importance measures to quantify component contributions to mixture effects. Using time series data from the National Morbidity, Mortality and Air Pollution Study, we demonstrate the proposed methods by jointly modelling three mortality outcomes and two cumulative air pollution measurements with a maximum lag of 14 days.
@@ 1 Introduction
Quantifying associations between health outcomes and mixtures of environmental pollutants is a public health priority, and addressing the methodological challenges introduced by mixtures analyses has been the focus of much recent research (Davalos et al.,, 2017; Stafoggia et al.,, 2017; Hamra and Buckley,, 2018; Maitre et al.,, 2022; Yu et al.,, 2022; Tanner et al.,, 2020; Joubert et al.,, 2022). Common to mixtures studies is a lack of power—owing in part to limited sample sizes, highly correlated exposures, and small effect sizes. This is exacerbated by the fact that environmental exposures may have non-linear relationships with health outcomes, so the functional form of these relationships must also be estimated from limited data. Moreover, in studies of time-varying exposures, the timing of exposure effects is itself unknown: for example studies of prenatal exposure to metals attempt to identify windows of time in which children are most susceptible to exposure effects (Barr Jr et al.,, 2000).
@@
Many methods have been proposed to solve one or more of these issues. To handle high correlation among mixture components in a linear model, Dirichlet Process (DP) priors have been placed on coefficients (Herring,, 2010; Dunson et al.,, 2008; LaLonde et al.,, 2020), among other latent variable approaches (Zavez et al.,, 2023). Non-linear relationships have been captured by semi-parametric regression using splines (Antonelli et al.,, 2020; Bai et al.,, 2022) and Gaussian processes (Bobb et al.,, 2015; Ferrari and Dunson,, 2019), as well as by single (Wang et al.,, 2020) and multiple index models (McGee et al.,, 2021, 2024). When exposures are measured repeatedly over time, distributed lag models (DLMs) have smoothed association parameters over time and characterized time-lag profiles that identify periods in which risks are highest (Gasparrini et al.,, 2010; Gasparrini,, 2014; Gasparrini et al.,, 2017; Wilson et al.,, 2017, 2022; Antonelli et al.,, 2024). While most of these approaches are intended to address multiple exposures, multiple time points, or both, none additionally address the complexity stemming from having multiple outcomes. It is common in environmental research to be interested in a variety of health outcomes, and analyzing each outcome separately ignores the fact that associations might be similar across outcomes.
@@
In this paper we propose a general approach to analyzing environmental mixtures that applies in single time point or time-varying exposure settings, which additionally borrows information across multiple outcomes to improve statistical power. In the time-varying exposure setting, we present a novel multivariate distributed lag non-linear model (MV-DLNM) and propose a strategy for borrowing strength across all exposure-outcome pairs by incorporating priors that induce clustering. In particular, we propose a co-clustering prior distribution that encourages similarity in both exposure-response curves as well as in time-lag profiles. Outside the time-varying exposure setting, the method applies immediately to other exposure structures (like exposures measured via multiple biomarkers) as well as settings with entirely unstructured exposures—for example borrowing strength across exposures in additive models or across outcomes in single index models. We then extend the proposed approach to fitting multiple index models (MIMs): with a minor modification, we show how the method can be used to fit MIMs when the true index structure—the number of indices and their composition—is unknown a priori. MIMs without known index groupings can be difficult to interpret, however, and we propose exposure importance measures that quantify the strength of each component’s contribution to the mixture.
2 Problem setup and modeling framework
For the $i^{th}$ study unit ($i=1,\dots,n$), suppose we observe a $K$-dimensional outcome vector $\mathbf{y}_{i}=(y_{i1},\dots,y_{iK})^{T}$, $P$ exposures at $L$ times: ${\mathbf{x}}_{i1},\dots,{\mathbf{x}}_{iP}$ where ${\mathbf{x}}_{ij}=({x}_{ij1},\cdots,{x}_{ijL})^{T}$, as well as a vector of other covariates $\bm{z}_{i}$. The goal is to quantify possibly non-linear associations between each outcome $y_{ik}$ and each time-varying exposure $\bm{x}_{ij}$ by fitting the following multivariate distributed lag non-linear model (MV-DLNM):
@@ 2 Problem setup and modeling framework
$y_{i1}$ $=f_{11}({\mathbf{x}}_{i1}^{T}{\bm{\omega}}_{11})+\cdots+f_{1P}({\mathbf{x}}_{iP}^{T}{\bm{\omega}}_{1P})+\bm{z}_{i}^{T}\bm{\beta}_{Z1}+\xi\sigma_{1}u_{i}+\epsilon_{i1}$ (1)
[…]
$y_{iK}$ $=f_{K1}({\mathbf{x}}_{i1}^{T}{\bm{\omega}}_{K1})+\cdots+f_{kj}({\mathbf{x}}_{iP}^{T}{\bm{\omega}}_{kj})+\bm{z}_{i}^{T}\bm{\beta}_{ZK}+\xi\sigma_{k}u_{i}+\epsilon_{iK}.$
@@
Here $f_{kj}(\cdot)$ is an unknown smooth exposure response curve for outcome $k$ and exposure $j$, ${\bm{\omega}}_{kj}=({\omega}_{kj1},\dots,{\omega}_{kjL})^{T}$ is a vector of unknown time-lag weights, and $\bm{\beta}_{Zk}$ is a vector of coefficients corresponding to covariates $\bm{z}_{i}$. We let $\epsilon_{ik}$ represent independent Gaussian errors, $\epsilon_{ik}\overset{iid}{\sim}N(0,\sigma^{2})$ and allow for dependence across the same unit’s outcomes via shared random effect $u_{i}\overset{iid}{\sim}N(0,1)$, scaled by the coefficient $\xi$. To identify both $f_{kj}(\cdot)$ and $\bm{\omega}_{kj}$, constraints on the weights’ magnitude (${\bm{\omega}}_{kj}^{T}{\bm{\omega}}_{kj}=1$) and sign (e.g., ${{\omega}}_{kj1}>0$) are typically imposed.
2.1 Semiparametric modeling incorporating smoothness
The target parameters of interest are the exposure-response function $f_{kj}(\cdot)$ and the lag weights ${\bm{\omega}}_{k}=({\omega}_{k1},\dots,{\omega}_{kL})^{T}$. The exposure-response function quantifies the potentially non-linear association between the $k^{th}$ outcome and the $j^{th}$ exposure, while the lag weights quantify the relative strength of an exposure’s association at each of the $L$ time periods. Importantly, the model specification allows each of the $K\times P$ outcome-exposure pairs to have (i) different non-linear exposure-response relationships $f_{kj}(\cdot)$, and (ii) different lag profiles captured by ${\bm{\omega}}_{kj}$. To facilitate estimation and improve efficiency, we incorporate smoothness conditions on both of the target parameters separately.
Smoothness over time: Exposure measurements over time are typically highly multicollinear, making estimation of the weights $\bm{\omega}_{kj}$ challenging. To improve stability, we can encourage smoothness in the weights $\bm{\omega}_{kj}$ over time via Gaussian priors with precision matrix $\lambda^{\omega}\mathbf{D}_{r}^{T}\mathbf{D}_{r}$, where $\mathbf{D}_{r}$ is a known difference matrix of order $r$ (Gasparrini et al.,, 2017). A prior with this covariance structure effectively penalizes or shrinks differences between coefficients at similar lags, enforcing our prior belief that effects should not change drastically over small time periods. The degree of shrinkage is driven by $\lambda^{\omega}$, an unknown parameter to be estimated.
While this prior encourages smoothness over time, it leaves the magnitude of ${\bm{\omega}}_{kj}$ completely unconstrained. To address this, we propose an analogous Fisher-Bingham prior (Kent,, 1982), which corresponds to a Gaussian distribution constrained to the unit $L$-sphere, thereby enforcing the constraint ${\bm{\omega}}_{kj}^{T}{\bm{\omega}}_{kj}=1$. Incorporating the same smoothness inducing covariance structure yields the prior $\bm{\omega}_{kj}\sim FB(\mathbf{0},\lambda^{\omega}\mathbf{D}_{r}^{T}\mathbf{D}_{r})$. This addresses the magnitude constraint required for $\bm{\omega}_{kj}$. To further satisfy the sign constraint, we can restrict this prior to the half hypersphere where $\omega_{kjL}\geq 0$. The constraints on the parameter space complicate Markov chain Monte Carlo (MCMC) sampling; in Section 4 we outline two approaches to sampling these parameters.
Sampling $\bm{\omega}_{kj}$ can be computationally burdensome when $L$ is large, however. To facilitate sampling, we can incorporate a dimension reduction approach similar to the one explored for DLMs in Wilson et al., (2022). We approximate ${\bm{\omega}}_{kj}\approx\mathbf{\Psi}{\bm{\theta}}_{kj}$, where $\mathbf{\Psi}$ is a (known) matrix whose columns define an orthonormal basis expansion, and ${\bm{\theta}}_{kj}$ is an $m$-dimensional vector of unknown coefficients, where ${\bm{\theta}}_{kj}$ inherits the magnitude constraint (${\bm{\theta}}_{kj}^{T}{\bm{\theta}}_{kj}=1$) by orthonormality. Smaller values of $m$ ($m\leq L$) lead to worse approximations, but easier posterior sampling. The covariance matrix that shrinks towards smoother functions is inherited as well, yielding the proposed prior $\bm{\theta}_{kj}\sim FB(\mathbf{0},\lambda^{\omega}\mathbf{\Psi}^{T}\mathbf{D}_{r}^{T}\mathbf{D}_{r}\mathbf{\Psi})$, and we can again impose a sign constraint ${{\theta}}_{kjL}>0$, restricting to the half-sphere. In sum, we apply a known linear transformation to ${\mathbf{x}}_{ij}$ and re-parameterize ${\mathbf{x}}_{ij}^{T}{\bm{\omega}}_{kj}\approx\tilde{\bm{x}}_{ij}^{T}{\bm{\theta}}_{kj}$. This dimension reduction is not required, and we could instead perform inference directly on $\bm{\omega}_{kj}$. This approximation simply serves to improve sampling; in Section 5 we explore the degree to which it affects inference. Additionally, if ${\mathbf{x}}_{ij}$ does not represent an exposure measured over time, but instead some other structure—say, an exposure measured via several biomarkers—we simply set $\mathbf{x}_{ij}=\tilde{\mathbf{x}}_{ij}$ and ${\bm{\omega}}_{kj}={\bm{\theta}}_{kj}$, and adopt a generic ridge penalty prior on ${\bm{\omega}}_{kj}$.
Smoothness of exposure-response functions: We encourage smoothness of $f_{kj}(\cdot)$ as well, as we expect small changes in exposure levels to lead to small changes in outcomes. We approximate $f_{kj}(\cdot)$ via B-splines: $f_{kj}(\tilde{{\mathbf{x}}}_{ij}^{T}{\bm{\theta}}_{kj})\approx\mathbf{b}_{ij,\bm{\theta}_{kj}}^{T}\bm{\beta}_{kj}$ where $\mathbf{b}_{ij,\bm{\theta}_{kj}}=(b_{1}(\tilde{\mathbf{x}}_{ij}^{T}\bm{\theta}_{kj}),\dots,b_{d}(\tilde{\mathbf{x}}_{ij}^{T}\bm{\theta}_{kj}))^{T}$, $b_{1}(\cdot),\dots,b_{d}(\cdot)$ are known basis functions, and $\bm{\beta}_{kj}=(\beta_{kj1},\dots,\beta_{kjd})^{T}$ is a vector of unknown coefficients. For identifiability, these basis representations are subject to centering constraints (Wood,, 2017), and we include outcome-specific intercepts $\beta_{k0}$, $k=1,\dots,K$. To encourage smoothness, we specify Gaussian priors on $\bm{\beta}_{kj}$ with precision matrix $\lambda^{\beta}\bm{\Sigma}^{-1}_{0}$, where $\bm{\Sigma}_{0}$ is a matrix of integrated second derivatives of basis functions that penalize overly rough functions (Wood,, 2017), and $\lambda^{\beta}$ is a smoothing parameter to be estimated.
2.2 Prior distributions inducing clustering
@@
The proposed model allows all $K\times P$ exposure-outcome pairs to have distinct time-lag profiles ($\bm{\omega}_{kj}\neq\bm{\omega}_{k^{\prime}j^{\prime}}$) and separate functional relationships ($\bm{\beta}_{kj}\neq\bm{\beta}_{k^{\prime}j^{\prime}}$). In small samples, or when the correlation among exposures across time is high, as is common in environmental epidemiology, this may be too flexible and lead to highly variable estimates. This can negatively affect power, which is already limited in the common situation where exposure effects are weak. To rein in this flexibility, we propose to encourage similarity in both $f_{kj}(\cdot)$ and in ${\bm{\omega}}_{kj}$ across different outcomes and exposures. This is a reasonable assumption in practice as many exposures have similar exposure-response curves, and the same exposure may have similar associations with different outcomes.
@@
A natural way to borrow information would be to adopt independent DP priors. Specifically, we can set $\bm{\beta}_{kj}\sim G_{\beta}$ where $G_{\beta}\sim DP(\alpha_{\beta},G_{0,\beta})$, and $\bm{\theta}_{kj}\sim G_{\theta}$ where $G_{\theta}\sim DP(\alpha_{\theta},G_{0,\theta})$. Here $(\alpha_{\beta},\alpha_{\theta})$ are concentration parameters, dictating the degree of clustering, and $(G_{0,\beta},G_{0,\theta})$ are centering distributions, representing the means of the prior distributions. To see how this encourages clustering and borrowing of information across exposure-outcome pairs, consider the following latent variable representation (Sethuraman,, 1994):
@@
$P(Z^{\beta}_{kj}=c)$ $=\pi_{c}^{\beta}$ $P(Z^{\theta}_{kj}=c)$ $=\pi_{c}^{\theta}~{}~{}c=1,2,\dots$
$\pi^{\beta}_{c}$ $=V^{\beta}_{c}\prod_{j<c}(1-V^{\beta}_{j})$ $\pi^{\theta}_{c}$ $=V^{\theta}_{c}\prod_{j<c}(1-V^{\theta}_{j})~{}~{}c=2,3,\dots,$
@@
Here $(Z^{\beta}_{kj},Z^{\theta}_{kj})$ is a discrete latent variable that indicates which cluster the exposure-outcome pair belongs for the exposure-response function and lag weights, respectively. In practice we approximate the infinite mixture with a finite mixture that is capped at $C$ clusters ($V^{\beta}_{C}=V^{\theta}_{C}=1$). It is easy to show that the prior probabilities of clustering are given by $P(\bm{\beta}_{kj}=\bm{\beta}_{k^{\prime}j^{\prime}})=(\alpha_{\beta}+1)^{-1}$ and $P(\bm{\theta}_{kj}=\bm{\theta}_{k^{\prime}j^{\prime}})=(\alpha_{\theta}+1)^{-1}$. This implies that clustering will occur for both parameters, and becomes increasingly likely with smaller values of $(\alpha_{\beta},\alpha_{\theta})$.
@@
Although this encourages clustering separately for $\bm{\beta}_{kj}$ and $\bm{\theta}_{kj}$, it does not encourage borrowing of information between the two. In practice, exposures may behave similarly both in terms of the shape of the exposure-response relationship as well as time-lag profile. For example two pollutants may both have U-shaped effects that decay linearly over the course of a week. Further, different outcomes may be susceptible to exposures at similar times and in similar manners. With that in mind we extend the DP clustering strategy above to encourage borrowing information across the exposure-response function and the time-lag domains. Consider the same latent variable representation as above. Instead of independently specifying marginal probabilities of clustering indicators $Z_{kj}^{\beta}$ and $Z_{kj}^{\theta}$, we specify a joint probability distribution that allows for dependence between the two:
$P(Z_{kj}^{\beta}=a,Z_{kj}^{\theta}=b)$ $=\frac{(1+\rho)^{I(a=b)}{\pi}^{\beta}_{a}{\pi}^{\theta}_{b}}{\sum_{a^{\prime}=1}^{C}\sum_{b^{\prime}=1}^{C}(1+\rho)^{I(a^{\prime}=b^{\prime})}{\pi}^{\beta}_{a^{\prime}}{\pi}^{\theta}_{b^{\prime}}},~{}~{}a,b\in\{1,\dots,C\}$

### v2

# Borrowing Strength Across Exposures and Outcomes via Index Models for Multi-pollutant Mixtures

@@ Abstract
An important goal of environmental health research is to assess the health risks posed by mixtures of multiple environmental exposures. In these mixtures analyses, flexible models like Bayesian kernel machine regression and multiple index models are appealing because they allow for arbitrary non-linear exposure-outcome relationships. However, this flexibility comes at the cost of low power, particularly when exposures are highly correlated and the health effects are weak, as is typical in environmental health studies. We propose a multivariate index modelling strategy that borrows strength across exposures and outcomes by exploiting similar mixture component weights and exposure-response relationships. In the special case of distributed lag models, in which exposures are measured repeatedly over time, we jointly encourage co-clustering of lag profiles and exposure-response curves to more efficiently identify critical windows of vulnerability and characterize important exposure effects. We then extend the proposed approach to the multiple index model setting where the true index structure—the number of indices and their composition—is unknown, and introduce variable importance measures to quantify component contributions to mixture effects. Using time series data from the National Morbidity, Mortality and Air Pollution Study, we demonstrate the proposed methods by jointly modelling three mortality outcomes and two cumulative air pollution measurements with a maximum lag of 14 days.
@@ 1 Introduction
Quantifying associations between health outcomes and mixtures of environmental pollutants is a public health priority, and addressing the methodological challenges of mixtures analyses has been the focus of much recent research (davalos2017current; stafoggia2017statistical; hamra2018environmental; maitre2022state; yu2022review; tanner2020environmental; joubert2022powering). Common to mixtures studies is a lack of power—owing in part to limited sample sizes, highly correlated exposures, and small effect sizes. This is exacerbated by the fact that environmental exposures may have non-linear relationships with health outcomes, so the functional form of these relationships must also be estimated from limited data. Moreover, in studies of time-varying exposures, the timing of exposure effects is itself unknown: for example, studies of prenatal exposure to metals attempt to identify windows of time in which children are most susceptible to exposure effects (barr2000workshop).
@@
Many methods address one or more of these issues. To handle high correlation among mixture components in a linear model, Dirichlet Process (DP) priors have been placed on coefficients (herring2010nonparametric; dunson2008bayesian; lalonde2020discovering), among other latent variable approaches (zavez2023bayesian). Non-linear relationships have been captured using splines (antonelli2020estimating; bai2022spike) and Gaussian processes (bobb2015bayesian; ferrari2019identifying), as well as by single (wang2020singleindex) and multiple index models (mcgee2021bayesian; mcgee2024collapsible). When exposures are measured repeatedly over time, distributed lag models have smoothed associations over time and characterized lag profiles to identify high risk time windows (gasparrini2010distributed; gasparrini2014modeling; gasparrini2017penalized; wilson2017bayesian; wilson2022kernel; antonelli2024multiple). While most of these approaches are intended to address multiple exposures, multiple time points, or both, few additionally address the complexity stemming from having multiple outcomes (warren2012bayesian; jain2018multivariate).
It is common in environmental research to be interested in multiple health outcomes, but analyzing each separately ignores the fact that associations might be similar across outcomes. For example, in the motivating analysis of cause-specific mortality and air pollution in Chicago (1987–2000), separately modeling lagged effects for each outcome yields wide intervals for associations between pollutants and respiratory/other non-accidental deaths, which could be compatible with either highly positive or negative associations.
@@
In this paper we propose a flexible index modelling framework for analyzing the relationships between environmental mixtures and health outcomes, and develop a co-clustering prior structure that borrows information across multiple outcomes and indices to improve power. The framework contains several important special cases, including a novel multivariate distributed lag non-linear model (MV-DLNM) when exposures are measured over time, in which the proposed co-clustering prior exploits similarity in exposure-response curves as well as in time-lag profiles. In our application to the Chicago air pollution mortality analysis, the proposed approach exploits similarity across several associations, resulting in more efficient inference and evidence of null relationships, while still capturing the strong non-linear effect of ozone on cardiovascular mortality. The framework also extends naturally to other exposure structures, such as exposures measured via multiple biomarkers. When exposures are measured only once, we show how the method can be used to fit multiple index models (MIMs) when the true index structure—the number of indices and their composition—is unknown a priori. To aid interpretability, we also propose exposure importance measures that quantify the strength of each component’s contribution to the mixture.
2 Methods
2.1 General Model Formulation and Parameterization
For the $i^{th}$ study unit ($i=1,\dots,n$), let $\mathbf{y}_{i}=(y_{i1},\dots,y_{iK})^{T}$ be a $K$-dimensional outcome vector, and let $\mathbf{x}_{i}^{*}$ be a general exposure vector of length $M$. Examples include: (i) $P$ exposures measured at $L$ times ($M=PL)$; (ii) $P$ exposures measured at $B$ different biomarkers ($M=PB$); (iii) $P$ exposures measured cross-sectionally ($M=P$). Finally let $\bm{z}_{i}$ be a vector of other covariates. For a given $J$, we write the proposed model:
@@ 2.1 General Model Formulation and Parameterization
$y_{ik}$ $=\sum_{j=1}^{J}f_{kj}\left([\mathbf{A}_{j}\mathbf{x}^{*}_{i}]^{T}{\bm{\omega}}_{kj}\right)+\bm{z}_{i}^{T}\bm{\beta}_{Zk}+\xi\sigma_{k}u_{i}+\epsilon_{ik},~~k=1,\dots,K,$ (1)
@@
where $\mathbf{A}_{j}$ is an $r\times M$ design matrix chosen by the analyst. The choice of $\mathbf{A}_{j}$ determines the model structure; we highlight two special cases in Sections 3 and 4, and outline others in Appendix A. Here $f_{kj}(\cdot)$ is an unknown smooth exposure-response function, $\bm{\omega}_{kj}$ is an $r$-vector of unknown weights, and $\bm{\beta}_{Zk}$ is a vector of coefficients corresponding to covariates $\bm{z}_{i}$. We let $\epsilon_{ik}$ represent independent Gaussian errors, $\epsilon_{ik}\overset{iid}{\sim}N(0,\sigma_{k}^{2})$ and allow for dependence across the same unit’s outcomes via shared random effect $u_{i}\overset{iid}{\sim}N(0,1)$, scaled by the coefficient $\xi$. We adopt a random intercepts structure here for simplicity, but more flexible random effect structures could be included with minor modification to the estimation procedure. To identify both $f_{kj}(\cdot)$ and $\bm{\omega}_{kj}$, constraints on the weights’ magnitude (${\bm{\omega}}_{kj}^{T}{\bm{\omega}}_{kj}=1$) and sign (e.g., ${{\omega}}_{kj1}>0$) are typically imposed.
For generality, we approximate ${\bm{\omega}}_{kj}\approx\mathbf{\Psi}{\bm{\theta}}_{kj}$, where $\mathbf{\Psi}$ is a known $r\times m$ matrix whose columns define an orthonormal basis as in wilson2022kernel, and ${\bm{\theta}}_{kj}$ is an $m$-dimensional ($m\leq r$) vector of unknown coefficients. Equivalently, $\tilde{\mathbf{x}}_{ij}^{T}{\bm{\theta}}_{kj}\approx[\mathbf{A}_{j}\mathbf{x}_{i}^{*}]^{T}\bm{\omega}_{kj}$, where $\tilde{\mathbf{x}}_{ij}=\mathbf{\Psi}^{T}\mathbf{A}_{j}\mathbf{x}_{i}^{*}$. Typically we avoid the approximation and set $\bm{\omega}_{kj}=\bm{\theta}_{kj}$, i.e. $\mathbf{\Psi}=\mathbf{I}_{r\times r}$, but we allow for $\mathbf{\Psi}$ with $m<r$ to facilitate dimension reduction and improved computation when $r$ is large. We approximate $f_{kj}(\cdot)$ via B-splines: $f_{kj}(\tilde{{\mathbf{x}}}_{ij}^{T}{\bm{\theta}}_{kj})\approx\mathbf{b}_{ij,\bm{\theta}_{kj}}^{T}\bm{\beta}_{kj}$ where $\mathbf{b}_{ij,\bm{\theta}_{kj}}=[b_{1}(\tilde{\mathbf{x}}_{ij}^{T}\bm{\theta}_{kj}),\dots,b_{d}(\tilde{\mathbf{x}}_{ij}^{T}\bm{\theta}_{kj})]^{T}$, $b_{1}(\cdot),\dots,b_{d}(\cdot)$ are known basis functions, and $\bm{\beta}_{kj}=(\beta_{kj1},\dots,\beta_{kjd})^{T}$ is a vector of unknown coefficients. For identifiability, the basis is subject to centering constraints (wood2017generalized), and we include outcome-specific intercepts $\beta_{k0}$.
2.2 Cluster-Inducing Prior Structure
@@
In the proposed model, there are $K\times J$ distinct functions $f_{kj}(\cdot)$ and $K\times J$ distinct weight profiles $\bm{\omega}_{kj}$, which can be overly flexible. In small samples, or when the correlation across indices is high, as is common in environmental epidemiology, this may lead to instability and low power, which is already limited in environmental studies prone to weak effect sizes. To rein in this flexibility, we exploit similarity in both $f_{kj}(\cdot)$ and in ${\bm{\omega}}_{kj}$ across pairs $(k,j)$. For example when $\mathbf{A}_{j}\mathbf{x}_{i}^{*}$ represents the $j^{th}$ exposure measured at $L$ different times (as in Section 3), this matches our expectation that exposures have similar exposure-response curves, and the same exposure may have similar associations with different outcomes.
@@
A natural way to borrow information would be to adopt independent DP priors: $\bm{\beta}_{kj}\sim G_{\beta}$ where $G_{\beta}\sim DP(\alpha_{\beta},G_{0,\beta})$, and $\bm{\theta}_{kj}\sim G_{\theta}$ where $G_{\theta}\sim DP(\alpha_{\theta},G_{0,\theta})$. Here $(\alpha_{\beta},\alpha_{\theta})$ are concentration parameters, dictating the degree of clustering, and $(G_{0,\beta},G_{0,\theta})$ are centering distributions. To see how this encourages clustering across exposure-outcome pairs, consider the following latent variable representation (sethuraman1994constructive):
@@
$P(Z^{\beta}_{kj}=c)$ $=\pi_{c}^{\beta}=V^{\beta}_{c}\prod_{j<c}(1-V^{\beta}_{j})$ $P(Z^{\theta}_{kj}=c)$ $=\pi_{c}^{\theta}=V^{\theta}_{c}\prod_{j<c}(1-V^{\theta}_{j})~~c=1,2,\dots$
@@
Here $(Z^{\beta}_{kj},Z^{\theta}_{kj})$ is a discrete latent variable that indicates to which cluster the pair $(k,j)$ belongs for $\bm{\beta}_{kj}$ and $\bm{\theta}_{kj}$, respectively. In practice we use a finite mixture with $C$ clusters ($V^{\beta}_{C}=V^{\theta}_{C}=1$), as there are only a finite number of parameters to be estimated, since there are $J\times K$ parameter vectors. Therefore we recommend setting $C=J\times K$ or to some fraction of this in order to speed up computation time without losing model flexibility. We assess sensitivity to this choice in Appendix G.4 and find that results are not sensitive to this choice as long as $C$ is large enough. It is easy to show that the prior probabilities of clustering are given by $P(\bm{\beta}_{kj}=\bm{\beta}_{k^{\prime}j^{\prime}})=(\alpha_{\beta}+1)^{-1}$ and $P(\bm{\theta}_{kj}=\bm{\theta}_{k^{\prime}j^{\prime}})=(\alpha_{\theta}+1)^{-1}$, and is increasingly likely with smaller values of $(\alpha_{\beta},\alpha_{\theta})$.
@@
Although this encourages clustering separately for $\bm{\beta}_{kj}$ and $\bm{\theta}_{kj}$, it does not encourage borrowing information between the two. In practice, exposures may behave similarly both in terms of the shape of the exposure-response relationship as well as weight profile. For example in DLNMs (where $(k,j)$ corresponds to an outcome-exposure pair, see Section 3), two pollutants may both have U-shaped effects that decay linearly over the course of a week, or different outcomes may be susceptible to exposures at similar times and in similar manners. With that in mind we extend the DP clustering strategy above to encourage borrowing information across the $\bm{\beta}_{kj}$ and $\bm{\theta}_{kj}$ domains. Consider the latent variable representation as above. Instead of independently specifying marginal probabilities of clustering indicators $Z_{kj}^{\beta}$ and $Z_{kj}^{\theta}$, we specify a joint distribution that allows for dependence between the two:
$P(Z_{kj}^{\beta}=a,Z_{kj}^{\theta}=b)$ $=\frac{(1+\rho)^{I(a=b)}{\pi}^{\beta}_{a}{\pi}^{\theta}_{b}}{\sum_{a^{\prime}=1}^{C}\sum_{b^{\prime}=1}^{C}(1+\rho)^{I(a^{\prime}=b^{\prime})}{\pi}^{\beta}_{a^{\prime}}{\pi}^{\theta}_{b^{\prime}}},~~a,b\in\{1,\dots,C\}.$